Assuming that the wave speed on a stretch string depends on the linear mass density meu as v inversely proportional to f/mue use dimentianal analysis to show that a =1/2and b=1/2
1
Expert's answer
2020-12-28T08:36:00-0500

Dimension of wave speed=[LT-1]

Dimension of force=[MLT-2]

Dimension of linear mass density=[ML-1].

We have

"v=\frac{f^a}{\mu^b}"

Where v is called wave speed ,f is called force and "\mu" is called linear mass density.Now putting dimension of each quantity in above equation

"[LT^{-1}]=\frac{[MLT^{-2}]^a}{[ML^{-1}]^b}"


"[LT^{-1}] =[M^{a-b}L^{a+b}T^{-2a}]"

On comparing both side we get

"-2a=-1"

"a=\frac{1}{2}"

And "a-b=0"

"a=b"

"b=\frac{1}{2}" ("\because" "a=\frac{1}{2}" )